David Rivera Optical has determined that its reorder point for eyeglass frames is 50 (d x LT) units. Its carrying cost per frame per year is $5, and stockout (or lost sale) cost is $40 per frame. The store has experienced the following probability distribution for inventory demand during the reorder period. The optimum number of orders per year is six.How much safety stock should David Rivera keep on hand?

Units Prob.
30 0.2
40 0.2
50 (ROP) 0.3
60 0.2
70 0.1
Answer #1

Reorder point = 50

Total cost = Incremental Holding cost ( Safety stock * 5) + Stockout cost ( Expected stockouts * probability of stockout * Cost of stockout * Number of orders / year0

Total cost when ROP = 50, SS = 0

= 0 * 5 + (60-50)*0.2*40*6 + (70-50)*0.1*40*6= $960

Total cost when ROP = 60, SS = 10

= 10 * 5 + (70-60)*0.1*40*6= $290

Total cost when ROP = 70, SS = 20

= 20 * 5 + 0 = $100

Hence, 70 frames is the optimal order with a safety stock of 20.